# Why Machine Learning Models Systematically Underestimate Extreme Values II: How to Fix It with LatentNN

### v1

@@ VIII Conclusions
This effect, well known in the statistics literature for linear regression (Fuller1987), extends to neural networks. Neural networks trained on noisy inputs suffer from the same bias as linear regression, with coefficient slopes biased toward zero by a factor that scales with $\sigma_{x}/\sigma_{\rm range}$. Increasing model complexity does not mitigate the effect. This issue has received little attention in the machine learning literature because most applications involve high-SNR inputs where the effect is negligible. Astronomy is different, and as neural networks become increasingly central to astronomical inference, classical statistical considerations such as errors-in-variables become important.
@@
I thank David Weinberg for the encouragement that led to Paper I of this series. The original idea about errors-in-variables stemmed from my teaching of Astron 5550 (Advanced Data Analysis), during which I decided to discuss attenuation bias as explored in Paper I. While preparing for that lecture, I was inspired to read more into the errors-in-variables literature, also suggested by an anonymous referee during Paper I. Further discussion with Jeffrey Newman and Brett Andrews during a colloquium at UPitt inspired the extension of this work to deep learning, which ultimately led to this paper. I also thank Hans-Walter Rix for useful discussion. This research was supported by NSF Grant AST-2406729 and a Humboldt Research Award from the Alexander von Humboldt Foundation. Claude 4.5 was used for assistance with code development and manuscript copy-editing. The scientific content, analysis, and conclusions are the author’s, and all LLM-assisted changes were verified by the author.
@@ Appendix A Derivation of Attenuation Bias
$x_{\rm obs}$ $=x_{\rm true}+\delta_{x},$ (A1)
$y_{\rm obs}$ $=\beta x_{\rm true}+\delta_{y},$ (A2)
@@
${\rm Cov}(x_{\rm obs},y_{\rm obs})$ $={\rm Cov}(x_{\rm true}+\delta_{x},\beta x_{\rm true}+\delta_{y})$ (A3)
$=\beta\,{\rm Var}(x_{\rm true})=\beta\sigma_{\rm range}^{2},$ (A4)
@@
${\rm Var}(x_{\rm obs})={\rm Var}(x_{\rm true})+{\rm Var}(\delta_{x})=\sigma_{\rm range}^{2}+\sigma_{x}^{2}.$ (A5)
@@
$\mathbb{E}[\hat{\beta}]=\beta\frac{\sigma_{\rm range}^{2}}{\sigma_{\rm range}^{2}+\sigma_{x}^{2}}=\beta\cdot\lambda_{\beta},$ (A6)
@@ Appendix B Derivation of Deming Regression
$\mathcal{L}=\prod_{i=1}^{N}\mathcal{N}(x_{{\rm obs},i}|x_{{\rm true},i},\sigma_{x}^{2})\cdot\mathcal{N}(y_{{\rm obs},i}|\beta x_{{\rm true},i},\sigma_{y}^{2}).$ (B1)
@@
$E(\{x_{{\rm true},i}\},\beta)=\sum_{i=1}^{N}\left[\frac{(x_{{\rm obs},i}-x_{{\rm true},i})^{2}}{2\sigma_{x}^{2}}+\frac{(y_{{\rm obs},i}-\beta x_{{\rm true},i})^{2}}{2\sigma_{y}^{2}}\right].$ (B2)
@@
$\frac{\partial E}{\partial x_{{\rm true},i}}=-\frac{x_{{\rm obs},i}-x_{{\rm true},i}}{\sigma_{x}^{2}}-\frac{\beta(y_{{\rm obs},i}-\beta x_{{\rm true},i})}{\sigma_{y}^{2}}=0.$ (B3)
@@
$x_{{\rm true},i}=\frac{r\,x_{{\rm obs},i}+\beta y_{{\rm obs},i}}{r+\beta^{2}}.$ (B4)
@@
$\hat{\beta}=\frac{S_{yy}-rS_{xx}+\sqrt{(S_{yy}-rS_{xx})^{2}+4rS_{xy}^{2}}}{2S_{xy}}.$ (B7)
@@ Appendix C Extension to Nonlinear Functions
To demonstrate this, we test LatentNN on a cubic polynomial $f(x)=x^{3}$. We use a larger network capacity (2 layers $\times$ 256 hidden units) to ensure the network can represent the nonlinear function accurately. Figure 7 shows results at $\mathrm{SNR}_{x}=2$, a challenging regime where noise equals 50% of the signal range. The same weight decay hyperparameter search procedure from Section V is applied.

### v2

@@ 8 Conclusions
This effect, well known in the statistics literature for linear regression (Fuller, 1987), extends to neural networks. Neural networks trained on noisy inputs suffer from the same bias as linear regression, with coefficient slopes biased toward zero by a factor that scales with $\sigma_{x}/\sigma_{\rm range}$. Increasing model complexity does not mitigate the effect. This issue has received little attention in the machine learning literature because most applications involve high-SNR inputs where the effect is negligible. Astronomy is different, and as neural networks become increasingly central to astronomical inference, classical statistical considerations such as errors-in-variables become important.
@@
Acknowledgements.
I thank David Weinberg and Hans-Walter Rix for useful discussion and comments on the manuscript. The original idea about errors-in-variables stemmed from my teaching of Astron 5550 (Advanced Data Analysis), during which I decided to discuss attenuation bias as explored in Paper I. While preparing for that lecture, I was inspired to read more into the errors-in-variables literature, also suggested by an anonymous referee during Paper I. Further discussion with Jeffrey Newman and Brett Andrews during a colloquium at UPitt inspired the extension of this work to deep learning, which ultimately led to this paper. This research was supported by NSF Grant AST-2406729 and a Humboldt Research Award from the Alexander von Humboldt Foundation. Claude 4.5 was used for assistance with code development and manuscript copy-editing. The scientific content, analysis, and conclusions are the author’s, and all LLM-assisted changes were verified by the author.
@@ Appendix A Derivation of Attenuation Bias
$x_{\rm obs}$ $=x_{\rm true}+\delta_{x},$ (18)
$y_{\rm obs}$ $=\beta x_{\rm true}+\delta_{y},$ (19)
@@
${\rm Cov}(x_{\rm obs},y_{\rm obs})$ $={\rm Cov}(x_{\rm true}+\delta_{x},\beta x_{\rm true}+\delta_{y})$ (20)
$=\beta\,{\rm Var}(x_{\rm true})=\beta\sigma_{\rm range}^{2},$ (21)
@@
${\rm Var}(x_{\rm obs})={\rm Var}(x_{\rm true})+{\rm Var}(\delta_{x})=\sigma_{\rm range}^{2}+\sigma_{x}^{2}.$ (22)
@@
$\mathbb{E}[\hat{\beta}]=\beta\frac{\sigma_{\rm range}^{2}}{\sigma_{\rm range}^{2}+\sigma_{x}^{2}}=\beta\cdot\lambda_{\beta},$ (23)
@@ Appendix B Derivation of Deming Regression
$\mathcal{L}=\prod_{i=1}^{N}\mathcal{N}(x_{{\rm obs},i}|x_{{\rm true},i},\sigma_{x}^{2})\cdot\mathcal{N}(y_{{\rm obs},i}|\beta x_{{\rm true},i},\sigma_{y}^{2}).$ (24)
@@
$E(\{x_{{\rm true},i}\},\beta)=\sum_{i=1}^{N}\left[\frac{(x_{{\rm obs},i}-x_{{\rm true},i})^{2}}{2\sigma_{x}^{2}}+\frac{(y_{{\rm obs},i}-\beta x_{{\rm true},i})^{2}}{2\sigma_{y}^{2}}\right].$ (25)
@@
$\frac{\partial E}{\partial x_{{\rm true},i}}=-\frac{x_{{\rm obs},i}-x_{{\rm true},i}}{\sigma_{x}^{2}}-\frac{\beta(y_{{\rm obs},i}-\beta x_{{\rm true},i})}{\sigma_{y}^{2}}=0.$ (26)
@@
$x_{{\rm true},i}=\frac{r\,x_{{\rm obs},i}+\beta y_{{\rm obs},i}}{r+\beta^{2}}.$ (27)
@@
$\hat{\beta}=\frac{S_{yy}-rS_{xx}+\sqrt{(S_{yy}-rS_{xx})^{2}+4rS_{xy}^{2}}}{2S_{xy}}.$ (30)
@@ Appendix C Extension to Nonlinear Functions
To demonstrate this, we test LatentNN on a cubic polynomial $f(x)=x^{3}$. We use a larger network capacity (2 layers $\times$ 256 hidden units) to ensure the network can represent the nonlinear function accurately. Figure 7 shows results at $\mathrm{SNR}_{x}=2$, a challenging regime where noise equals 50% of the signal range. The same weight decay hyperparameter search procedure from Section 5 is applied.